Solve y′ = x−y^2, y(0) = 1 using the forward Euler method for in [0, 0.6] by taking h = 0.2.


1
Expert's answer
2021-02-01T08:28:14-0500

By Euler's formula we know that , "y_{n+1}=y_n+h.f(x_n,y_n)" .........(1)

Here given that "f(x,y)=x-y^2" with "y(0)=1."

"\therefore x_0=0" and "y_0=1"

Also given that "h=0.2" . We have to find the value of "y" in the interval "[0,0.6]" .

We also know that, "x_{n+1}=x_n+h"

Therefore, "x_1=0.2" for "n=0"

"x_2=0.4" for "n=1"

"x_3=0.6" for "n=2"

Now corresponding to these values of "x_1,x_2,x_3" we get the values of "y_1,y_2,y_3" with the help of (1).

For "n=0" , "y_1=y_0+h.f(x_0,y_0)"

"=1+(0.2)\u00d7[0-1^2]"

"=1-0.2=0.8"

For "n=1," "y_2=y_1+h.f(x_1,y_1)"

"=0.8+(0.2)\u00d7[0.2-(0.8)^2]"

"=(0.8-0.088)=0.712"

For "n=2," "y_3=y_2+h.f(x_2,y_2)"

"=0.712+(0.2)\u00d7[0.4-(0.712)^2]"

"=0.712-0.0213838"

"=0.6906"

Therefore the required solutions in "[0,0.6]" are , "y(0)=1" , "y(0.2)=0.8" , "y(0.4)=0.712" and "y(0.6)=0.6906" .